An article in the San Jose Mercury News stated that students in the California state university system take 4.5 years, on
average, to finish their undergraduate degrees. Suppose you believe that the mean time is longer. You conduct a survey of
49 students and obtain a sample mean of 5.1 with a sample standard deviation of 1.2. Do the data support your claim at the
1% level?

Answer by Boreal(15235)

Ho-mean is 4.5
Ha-mean is not 4.5
alpha=0.01
test is a t 0.995, df=48
t=(xbar-mu)/s/sqrt(n)
critical value=2.6822
reject Ho for |t|>2.6822
t=(4.5-5.1)/1.2/sqrt(49)=-4.2/1.2=-3.5
reject Ho, p-value=0.001
The claim is supported at the 1% level